OpenStudy (anonymous):

Find the distance between (-5, 3) and (-5, 9)

OpenStudy (anonymous):

distance formula

OpenStudy (rmrjr22):

c2 = (xA - xB)2 + (yA - yB)2

OpenStudy (anonymous):

distance formula is \[\sqrt{(x2-x1)^{2 +(y2-y1)^{2}}}\]

OpenStudy (anonymous):

0 squared plus 6 squared? i dont think im doing it right

OpenStudy (rmrjr22):

\[\sqrt{-10^2+6^2}\]

OpenStudy (anonymous):

\[\sqrt{-5-(-5)^{2} +(9-3)^{2}}\]

OpenStudy (rmrjr22):

\[\sqrt{100+36}\]

OpenStudy (rmrjr22):

i lied....

OpenStudy (rmrjr22):

just \[\sqrt{36}\]

OpenStudy (rmrjr22):

solve and u get distance

OpenStudy (anonymous):

6

OpenStudy (rmrjr22):

bingo

OpenStudy (rmrjr22):

Best Response? :)

OpenStudy (anonymous):

you got it lol

OpenStudy (rmrjr22):

thx

OpenStudy (anonymous):

I gave it to @heather040200

OpenStudy (anonymous):

Thanks @timo86m :)

OpenStudy (rmrjr22):

:(

OpenStudy (rmrjr22):

its cause shes pretty isnt it... :/

OpenStudy (anonymous):

i gave one to you @rmrjr22 :)

OpenStudy (rmrjr22):

Ditto

OpenStudy (anonymous):

Thanks were all just givin each other medals LOL

OpenStudy (anonymous):

She is pretty but she also got it right :D